Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.50 (given its target capital structure). Vandell has $8.23 million in debt that trades at par and pays an 7.2% interest rate. Vandell’s free cash flow (FCF0) is $2 million per year and is expected to grow at a constant rate of 4% a year. Both Vandell and Hastings pay a 40% combined federal and state tax rate. The risk-free rate of interest is 7% and the market risk premium is 4%. Hastings Corporation estimates that if it acquires Vandell Corporation, synergies will cause Vandell’s free cash flows to be $2.3 million, $3.0 million, $3.4 million, and $3.85 million at Years 1 through 4, respectively, after which the free cash flows will grow at a constant 4% rate. Hastings plans to assume Vandell’s $8.23 million in debt (which has an 7.2% interest rate) and raise additional debt financing at the time of the acquisition. Hastings estimates that interest payments will be $1.5 million each year for Years 1, 2, and 3. After Year 3, a target capital structure of 30% debt will be maintained. Interest at Year 4 will be $1.414 million, after which the interest and the tax shield will grow at 4%. Indicate the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition. Round your answers to the nearest cent. Do not round intermediate calculations. The bid for each share should range between $ per share and $ per share.
Answer #1

a) Vandells pre-acquisition levered cost of equity rSL = 7 + 4*1.5 = 13

rsU = 7.2*0.4+13*0.6 = 10.68= 10.68%   

Finding the WACC of Vandell Corporation, to dsicount the FCFs

Cost of Debt 7.2%*(1-40%)= 4.32%

WACC=(40%*4.32%)+(60%*10.68%)=8.136%

Vops=FCF0(1+g)/(WACC-g)=($2M(1+0.04))/(0.08136-0.04)=$27M

VS=Vops-debt=$27 M-$8.23 M=$18.77 M

Price=$18.77 M/1 M share=$18.77 per share

b) Intrinsic un levered value of operations (with synergy):

Year Cash Flow(Million)
1 $2.3
2 $3.0
3 $3.4
4 $3.85

b)

Tax Shields 1-3 = Interest * Tax Rate

Tax Shields 1-3 =$ 1.5M * 40%

Tax Shields 1-3 =$ 400,000

Tax Shield 4 = $1.414M * 40%=$565,600

Tax shield horizon value = TS4(1+g)/(rsU-g)

= 0.5656 (1.04)/(0.1068-0.04) = 8.80

Value of tax shields =(0.400/1.1068)+(0.400/(1.1068)2)+(0.400/(1.1068)3)+(0.5656+8.80/(1.1068)4)=$7.22 M

Unlevered horizon value = FCF4(1+g)/(rsU-g) = 3.85(1.04)/(0.1068-0.04) = 59.94

Unlevered Vops=(2.3/1.1068)+(3.0/(1.1068)2)+(3.4/(1.1068)3)+(3.85+59.94/(1.1068)4)=$49.53

Value of operations = unlevered Vops+ value of tax shields = 49.53+ 7.22 = $56.75 million

Equity value to Harrison = Vops – Debt =$ 56.75 million -$ 8.23 million = $ 48.52 M

or $48.52 per share since there are 1 million shares outstanding.

On the basis of the answers in a and b, the bid for each share should range between $18.77 and $48.52